### v1

# A Risk Taxonomy for Evaluating AI-Powered Psychotherapy Agents

Abstract.
The proliferation of Large Language Models (LLMs) and Intelligent Virtual Agents acting as psychotherapists presents significant opportunities for expanding mental healthcare access. However, their deployment has also been linked to serious adverse outcomes, including user harm and suicide, facilitated by a lack of standardized evaluation methodologies capable of capturing the nuanced risks of therapeutic interaction. Current evaluation techniques lack the sensitivity to detect subtle changes in patient cognition and behavior during therapy sessions that may lead to subsequent decompensation. We introduce a novel risk taxonomy specifically designed for the systematic evaluation of conversational AI psychotherapists. Developed through an iterative process including review of the psychotherapy risk literature, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria (e.g., DSM-5) and existing assessment tools (e.g., NEQ, UE-ATR), the taxonomy aims to provide a structured approach to identifying and assessing user/patient harms. We provide a high-level overview of this taxonomy, detailing its grounding, and discuss potential use cases. We discuss two use cases in detail: monitoring cognitive model-based risk factors during a counseling conversation to detect unsafe deviations, in both human-AI counseling sessions and in automated benchmarking of AI psychotherapists with simulated patients. The proposed taxonomy offers a foundational step towards establishing safer and more responsible innovation in the domain of AI-driven mental health support.
AI Psychotherapy, Risk Assessment, Intelligent Virtual Agents, LLM Safety, Mental Health Technology
journalyear: 2025conference: ; May 2025; booktitle: Preprintccs: Applied computing Consumer healthccs: Information systems Language modelsccs: Human-centered computing Empirical studies in HCI
1. Introduction
Artificial Intelligence (AI) systems, particularly those leveraging Large Language Models (LLMs), are increasingly being developed and deployed as tools for mental health support, including entertainment chatbots ”masquerading” as therapists (Jurblum and Selzer, 2025; Barry, 2025; Jin et al., 2025; Na et al., 2025). The American Psychological Association has recently warned federal regulators about AI chatbots programmed to reinforce, rather than challenge, a user’s thinking, potentially driving vulnerable users to harm themselves or others (Lawrence et al., 2024; Barry, 2025). Reports have emerged, including instances where interactions with AI chatbot ”therapists” have been linked to severe negative outcomes. Notably, one case involved a 14-year-old boy’s suicide after months of interaction with chatbots on Character.AI, including a chatbot that falsely claimed to have several years’ experience as a licensed therapist (Roose, 2024; Barry, 2025). In this instance, the chabot reportedly reinforced the user’s harmful ideations rather than challenging them (Barry, 2025). Such events underscore the profound ethical responsibilities and potential harms associated with deploying these technologies to vulnerable populations, often with concerns raised about them being ”dangerous and untested” or lacking rigorous, context-appropriate safety evaluations (Roose, 2024; Barry, 2025; Stade et al., 2024b).
A critical issue is the absence of standardized methods to evaluate these AI systems for the unique risks inherent in psychotherapeutic interactions. Existing LLM benchmarks often focus on general capabilities such as toxicity detection in isolated texts, failing to capture harms that may emerge from sustained, dynamic conversations (Chang et al., 2023; Zhang et al., 2025). Traditional clinical trials, while vital for efficacy, can be slow and ethically challenging for initial safety validation of unproven autonomous systems (Stade et al., 2024b; Heinz et al., 2025). This gap leaves users susceptible to unexpected adverse outcomes.
To address this, we propose a risk taxonomy tailored for evaluating AI psychotherapists. We present a high-level overview of this taxonomy, derived from an iterative process including a review of the literature on psychotherapy risks, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria and existing assessment tools. We hypothesize that this taxonomy can provide a structured and systematic approach to identifying, categorizing, and ultimately mitigating risks in AI-driven psychotherapy applications. Our primary aim is to provide a framework that can be used to evaluate automated psychotherapy systems for potential negative impacts before and during their deployment, fostering safer innovation in digital mental health.
2. Related Work
2.1. Risks in Psychotherapy
Psychotherapy, while generally effective, can occasionally lead to negative consequences (Linden and Schermuly-Haupt, 2014; Boisvert and Faust, 2002). Estimates of unwanted effects vary from 3% to over 50%, depending on the particular definitions and assessment methods used (Linden, 2013; Mejía-Castrejón et al., 2024). A key challenge is the lack of a uniform conceptual framework and standardized tools for assessing adverse events (AEs) in psychotherapy (Mejía-Castrejón et al., 2024).
Several frameworks attempt to categorize negative experiences in psychotherapy. Linden and Schermuly-Haupt distinguish Unwanted Events (UEs) from Adverse Treatment Reactions (ATRs) caused by correctly applied treatment, and malpractice reactions from incorrectly applied treatment (Linden, 2013; Linden and Schermuly-Haupt, 2014). Mejía-Castrejón et al. (Mejía-Castrejón et al., 2024) adapted a medical model, classifying AEs by severity, relatedness, seriousness, and expectedness. Common AEs identified across studies and captured in assessment tools like the Negative Effects Questionnaire (NEQ) (Rozental et al., 2019) or the Unwanted Event to Adverse Treatment Reaction (UE-ATR) checklist (Linden, 2013) include negative emotions (e.g., anxiety, sadness), symptom worsening, the emergence of unpleasant memories, relationship strains, dependency on therapy, or reduced self-efficacy (Linden and Schermuly-Haupt, 2014; Yao et al., 2020). Factors influencing these include patient and therapist characteristics, and specific techniques (Linden and Schermuly-Haupt, 2014).
Assessing these effects is difficult; therapists may struggle to recognize negative outcomes, and establishing causality is complex (Linden, 2013; Linden and Schermuly-Haupt, 2014). While tools like the UE-ATR, NEQ, and the Inventory for Negative Effects of Psychotherapy (INEP) (Ladwig et al., 2014) exist, consensus on definitions and instruments remains elusive (Mejía-Castrejón et al., 2024). Reviews show variability and inadequate reporting of negative effects in randomized clinical trials (Jonsson et al., 2014; Klatte et al., 2023), underscoring the need for standardized assessment. Our taxonomy development aims to build upon these existing efforts by providing a structured framework specially adapted for evaluating AI-powered psychotherapy agents.
2.2. Safety Evaluation of LLMs in Healthcare
The application of LLMs in high-stakes domains like healthcare, including psychotherapy, demands rigorous safety assessment due to the potential for significant harm (Stade et al., 2024b; Lawrence et al., 2024; Steenstra et al., 2024). General LLM safety evaluations often focus on risks including bias, toxicity, or misinformation (Weidinger et al., 2021; Ji et al., 2023), using benchmarks such as ALERT (Tedeschi et al., 2024) or SafetyBench (Zhang et al., 2023). However, these may not adequately capture the interactive and relational risks specific to psychotherapy.
Safety frameworks are emerging for AI in mental health, including the READI framework (Stade et al., 2024a) and the BPES checklist, emphasizing ”Do No Harm” (Khan and Seto, 2023). Dialogue safety in mental health support is also a focus (Qiu et al., 2023). For instance, $\Psi$-ARENA offers an interactive platform for evaluating LLM-based psychological counselors (Zhu et al., 2025). Despite these advances, a standardized taxonomy to systematically identify and categorize the spectrum of potential risks specific to AI psychotherapeutic interactions, grounded in clinical realities and existing harm assessment principles, is still lacking. This is crucial as AI systems, unlike human therapists, may lack nuanced understanding of human emotion, contextual subtleties, and the ability to dynamically repair relational ruptures, potentially leading to unique failure modes.
3. Taxonomy Development
The development of our risk taxonomy followed an iterative design process. We first formulated a preliminary taxonomy based on an in-depth literature review of psychotherapy risk (e.g., (Linden, 2013; Rozental et al., 2018; Klatte et al., 2023)). This initial framework subsequently underwent expert review through semi-structured interviews with 11 experts: 10 practicing psychologist professionals and one legal professional specializing in healthcare malpractice. These interviews were designed to explore perceptions of risk in human therapy, identify potential failure modes for AI therapists, and gather direct feedback on the preliminary taxonomy. The interview study was approved by Northeastern University’s IRB.
The interviews were analyzed using thematic analysis (Clarke and Braun, 2017) to reveal six key themes.
3.1. Theme 1: Therapy may require short-term discomfort
The first theme highlighted the difficulty in defining and identifying ”harm” within the therapeutic context itself. Unlike the clearer ”do no harm” principle in general medicine, psychotherapy often involves navigating ”intentional discomfort” as part of the therapeutic process. Participants described therapy inherently involving potentially negative feelings (”feeling a little worse before you feel better,” ”encounter some painful things”) as necessary for growth. Distinguishing this expected discomfort from ”unintentional harm”—negative outcomes resulting from therapist errors, misjudgments, or failures—is challenging. Harm exists on a ”spectrum,” ranging from temporary setbacks (”bruises”) like feeling ”misunderstood” or ”judged,” to more significant consequences (”cuts”) such as a ”rupture” in the ”therapeutic alliance,” premature termination (”stop coming to therapy,” ”drop out early”), worsening symptoms (”exacerbate symptoms”), loss of ”trust in self,” developing a ”bad taste… about therapy,” or, in the most severe cases, increased ”suicide” risk. Furthermore, harm can be subjective, context-dependent (”contextual based on your history”), and may not manifest ”in the moment,” only emerging later through ”rumination.”
3.2. Theme 2: Therapist risk factors
Therapist factors related to risk are significant, encompassing competence (”evidence-based care,” ”qualifications,” ”knowing their specialty”), ethical adherence (”violates boundaries,” ”confidentiality”), and skillful technique (avoiding ”leading questions,” ”invalidating” responses, ”negative vocabulary,” ”overly reassuring”). Therapist ”biases,” ”personal opinions,” lack of ”cultural awareness,” ”distractions,” and ”burnout” were cited as potential sources of unintentional harm. Building ”therapeutic rapport” and ”trust” was seen as a primary mitigator, allowing for ”repair” if harm occurs.
3.3. Theme 3: Patient risk factors
Patient vulnerability and context also play a critical role. ”Individual differences” mean risk factors vary; internal states (”hopelessness,” ”anxiety”), patient history (”memories,” ”interpretations”), the ”social environment,” ”physical environment” (including ”access to means”), and ”protective factors” all interact dynamically.
3.4. Theme 4: Real-time nature of risk assessment
Identifying risks in real-time relies heavily on a therapist’s ability for ”moment-to-moment tracking,” interpreting subtle ”nonverbal body language,” ”facial expressions,” and ”para verbal cues” (e.g., ”intonation”), alongside verbal content and formal ”assessment tools” (e.g., ”Columbia Scale”(Posner et al., 2008), ”PHQ-9”(Kroenke et al., 2001)). However, assessment is imperfect due to patient ”concealment,” the limits of prediction (”intention could change”), and the significant information loss in modalities lacking visual and auditory cues, such as ”texting” therapy, which was viewed as particularly risky by many participants. Critically, effective responses require understanding the ”underlying cause,” not just the presenting ”symptom.”
3.5. Theme 5: Perceived shortcomings of AI therapists
When considering AI therapists, participants largely believed the fundamental risks in human therapy remain the same but are potentially ”higher” or exacerbated. The primary concern stems from the AI’s inherent inability (especially in text-based interactions) to perceive the rich nonverbal and paralinguistic data humans use to gauge internal states, build rapport, and understand nuance (”missing the person-to-person component,” ”can’t synthesize,” loss of ”nuance”). This limitation raises concerns about AI’s capacity for genuine ”empathy,” its potential to ”misinterpret” user input, perpetuate ”bias,” lack crucial ”context,” miss the ”art” of therapy, and ultimately, its ability to safely navigate complex emotional terrain. While potential AI advantages like constant ”availability,” potentially feeling ”less judged,” and superior information processing were acknowledged, the loss of the human connection and observational capacity was seen as a major safety deficit. Interestingly, AI was often considered potentially ”better” than human texting therapy, as the latter shares AI’s lack of nonverbals without AI’s potential processing advantages.
3.6. Theme 6: Refinement of risk taxonomy
The feedback on the proposed risk taxonomy indicated it was a ”good start” covering many ”big risks,” but highlighted needs for refinement. Key omissions noted were ”trauma” and ”abuse.” Participants stressed the necessity of integrating ”environmental” and ”social factors,” the ”timeline” of symptoms, and grounding factors in established tools like the ”DSM” or clinical ”inventories,” while acknowledging the limitations of categorical approaches versus ”individual differences” and ”nuance.” The taxonomy’s current structure was seen as potentially useful but insufficient alone for risk assessment without incorporating deeper context and validated measures.
These insights drove significant revisions: patient factors within the risk taxonomy were linked to the DSM-5, factor values are now compared to baselines to accommodate individual differences, and risk categorization was grounded in approaches similar to the NEQ and UE-ATR.
4. Overview of the Risk Taxonomy
The refined risk taxonomy, derived from literature (Linden, 2013; Rozental et al., 2018; Klatte et al., 2023), expert interviews, and DSM-5 criteria (Black and Grant, 2014), provides a structured framework for identifying potential negative outcomes in AI psychotherapy. The full taxonomy can be found in Appendix A. It has two main categories:
1. Immediate Risk: This category encompasses situations of imminent danger to the user or others, requiring urgent intervention. Examples include explicit suicidal intent with plan and means, threats of harm to others, or severe psychological decompensation such as acute psychosis. Identification relies primarily on dialogue-based triggers. Evaluation assesses the AI’s ability to recognize these triggers and respond according to predefined, ethically-informed action plans, such as attempting de-escalation, recommending emergency resources (e.g., 911, 988 Lifeline), and contacting/flagging a human supervisor immediately.
2. Potential Risk: This category refers to emerging states or interaction patterns suggesting heightened vulnerability to negative therapeutic outcomes, functional deterioration, or adverse impacts on the therapeutic process, often aligning with unwanted events in the NEQ or UE-ATR (Rozental et al., 2019; Linden, 2013). Key subcategories are: 2.1. Symptom Exacerbation, Emergence of New Symptoms, or Maladaptive Coping, which involves worsening of clinical symptoms (e.g., increased negative thoughts or feelings) or new maladaptive behaviors (e.g., Worry/Rumination Frequency); 2.2. Patient Destabilization or Feeling Overwhelmed, indicating the user is struggling with the therapeutic process (e.g., increased confusion, emotional lability); 2.3. Therapeutic Alliance Damage or Negative Relational Dynamics, covering negative shifts in the user-AI relationship including mistrust, feeling judged, or over-dependence; and 2.4. Reduced Engagement or Potential for Premature Termination, signaling decreased participation or commitment within the therapeutic process or current therpay session.
Identification of these potential risks involves tracking dynamic changes in internal cognitive, emotional, and behavioral factors, which are conceptualized as scales (e.g., from 1-low to 10-high for intensity or frequency, such as ”Hopelessness Level” or ”Collaboration/Engagement Level”). The up (↑) and down (↓) arrows accompanying these factors in the full taxonomy (see Appendix A) specify the direction of change from a user’s baseline state that may signify an increased potential risk. For example, an increase (↑) in ”Hopelessness Level” or a decrease (↓) in ”Collaboration/Engagement Level” would be flagged as potentially problematic.
Evaluation should assess harmful shifts in these factors from baseline values, allowing for the tracking of subtle deteriorations or emerging vulnerabilities. While empirically-grounded numeric thresholds for what constitutes a significant change is a goal for future work, the current approach suggests that any major deviation beyond simple, minor fluctuations around a baseline cognitive factor value (e.g., more than +/- 1 point on a 1-10 scale from the baseline), or other abnormal fluctuations, could indicate an increased potential risk requiring attention. Furthermore, as expert participants highlighted, these indicators should ideally be monitored over the course of a session or across multiple interactions to understand their trajectory and contextual significance, rather than relying solely on isolated measurements. For instance, a consistent decrease (↓) in the Motivation/Self-Efficacy Level factor over multiple sessions could strongly suggest an increased risk for Reduced Engagement / Potential for Premature Termination.
5. Expected Use Cases and Adaptability
The taxonomy standardizes safety evaluation criteria for AI-powered psychotherapy agents. There are several ways the taxonomy can be used to improve safety. We outline several use cases here.
5.1. Monitoring Real User Interactions
The risk taxonomy could be used to identify emergent risks in real-time, as a user interacts with an AI therapist, for example to provide early termination of a harmful session, or feedback to developers or regulatory bodies. Monitoring should be performed with ethical consent, and could be accomplished by tracking user risk factors from pre- to post-interaction to identify ”Potential Risk” shifts, while dialogue tools scan for ”Immediate Risk” triggers.
5.2. Evaluation with Simulated Patients
The risk taxonomy could also be used to ”monitor” interactions between an AI therapist and a simulated patient, such as (Steenstra et al., 2025), providing a powerful evaluation testbed that can be used to assess systems for safety and potential harm. During these simulated interactions, simulated patient agents’ dynamic internal states corresponding to ”Potential Risk” factors (e.g., ”Hopelessness Level) are updated. This allows systematic testing by comparing baseline to post-interaction values of internal patient factors and observing AI responses to diverse personas, a methodology previously validated in our development of SimPatient (Steenstra et al., 2025), an LLM-powered system where embodied simulated patients featured dynamic internal cognitive states (e.g., self-efficacy) that evolved during interactions.
5.3. Benchmarking and Comparative Analysis
The risk taxonomy provides a standardized framework essential for robust benchmarking and comparative analysis of different AI psychotherapist models or system versions. By systematically evaluating AI therapists against the defined risk categories and factors, consistent and comparable risk profiles can be generated for each system. This process can highlight the relative strengths and weaknesses of various AI approaches in terms of their propensity to elicit or mitigate specific types of risks. For example, one model might excel at de-escalating immediate suicidal ideation but perform poorly by increasing therapeutic alliance damage, while another might show the opposite pattern. Such comparative insights are invaluable for developers seeking to enhance model safety, for organizations making informed deployment decisions based on specific use-case risk tolerances, and for fostering a more transparent and accountable development ecosystem focused on minimizing harm.
5.4. Identifying Unexpected Outcomes
Finally, the risk taxonomy serves as a crucial tool for identifying unexpected or emergent adverse outcomes that may arise during automated therapy. This is achieved through systematic monitoring of user interactions and therapeutic outcomes, continually assessed against the taxonomy’s defined risk factors and categories. Such monitoring can reveal significant deviations from established baseline values—which ideally exhibit only safe fluctuations—or flag the emergence of previously unrecognized harmful patterns. When substantial deviations or novel harmful patterns are detected, a deeper analysis is warranted. This involves scrutinizing the conversational context surrounding these occurrences to understand the specific interaction dynamics or AI behaviors that may have precipitated them. Uncovering these root causes is invaluable, as it directly informs targeted safety improvements and iterative refinements of the AI model, ultimately leading to a more robust and safer therapeutic tool.
While our risk taxonomy provides a general framework, it can also be augmented for specific populations or problems. For instance, evaluating an AI psychotherapist for helping patients with substance use disorders could add ”Potential Risk” factors such as ”Substance Craving Intensity”, ”Ambivalence about changing substance use” ”Substance Use Frequency/Amount,” ”Level of Intoxication/Withdrawal Symptoms”, and relational dynamics like ”Dishonesty/Minimization regarding substance use.” This tailoring allows for more nuanced risk assessment. Future work includes validating the taxonomy against human expert judgments.
6. Conclusion
AI psychotherapists present significant opportunities and risks. Documented negative outcomes demand standardized evaluations for unique interaction risks. We developed the risk taxonomy, from an iterative process involving literature, expert insights, and DSM-5, and harm checklists. It offers a structured approach to identifying and categorizing harms. Its overview, grounding, and adaptable use cases (tracking risk factor changes from baseline to post-interaction) aim to equip stakeholders to ensure AI psychotherapy safety, fostering innovation that prioritizes user well-being and minimizes unintended negative consequences.
Appendix A Risk Taxonomy